using the following balanced equation how many moles of water will be produced a 4.25 miles of oxygen is totally consumed. Show all work.

4NH3+3O2=2N2+6H20


1
Expert's answer
2022-04-07T03:36:06-0400

According to the equation, n (H2O) = 6/3 x n (O2) = 6/3 x 4.25 = 8.5 moles